OpenStudy (anonymous):

Determine if the series is absolutely convergent, conditionally convergent, or divergent. \[\sum_{2}^{\infty}n \div \left( \ln n \right)^{n}\]

OpenStudy (zarkon):

Absolutely convergent...use root test

OpenStudy (anonymous):

how do u use the root test, I'm still confused about it

OpenStudy (anonymous):

\begin{eqnarray*}\limsup_{n\to\infty}\left(\sqrt[n]{\frac{n}{\log^n(n)}}\right)&=&\limsup_{n\to\infty}\left(\frac{\sqrt[n]{n}}{\log(n)}\right)\\&=&\lim_{n\to\infty}\left(\frac{\sqrt[n]{n}}{\log(n)}\right)\\&=& 0.\end{eqnarray*}Because this number is less than 1, the series is absolutely convergent.